# CRUD-RAG: A Comprehensive Chinese Benchmark for Retrieval-Augmented Generation of Large Language Models

### v2

@@ Abstract.
Retrieval-Augmented Generation (RAG) is a technique that enhances the capabilities of large language models (LLMs) by incorporating external knowledge sources. This method addresses common LLM limitations, including outdated information and the tendency to produce inaccurate ”hallucinated” content. However, evaluating RAG systems is a challenge. Most current benchmarks mainly focus on question-answering applications, and neglect the wider range of scenarios where RAG could be beneficial. Moreover, they only evaluate the performance of LLM components of RAG pipeline in the experiments, and neglect the influence of retrieval components and external knowledge database construction. To address these issues, this paper constructs a large-scale and more comprehensive benchmark, and evaluates all the components of RAG systems in various RAG application scenarios. Specifically, we refer to the CRUD actions that describe interactions between users and knowledge bases, and also categorize the range of RAG applications into four distinct types—Create, Read, Update, and Delete (CRUD). ”Create” refers to scenarios requiring the generation of original, varied content. ”Read” involves responding to intricate questions in knowledge-intensive situations. ”Update” focuses on revising and rectifying inaccuracies or inconsistencies in pre-existing texts. ”Delete” pertains to the task of summarizing extensive texts into more concise forms. For each of these CRUD categories, we have developed different datasets to evaluate the performance of RAG systems. We also analyze the effects of various components of the RAG system, such as the retriever, context length, knowledge base construction, and LLM. Finally, we provide useful insights for optimizing the RAG technology for different scenarios [^1] The source code is available at GitHub: https://github.com/IAAR-Shanghai/CRUD_RAG.
@@ 1. Introduction
Although the effectiveness of retrieval-augmented strategies has been proven through extensive practice, their implementation still requires a significant amount of tuning. The overall performance of the RAG system is affected by multiple factors, such as the retrieval model, construction of the external knowledge base, and language model. Therefore, automatic evaluation of RAG systems is crucial. Currently, there are only a few existing benchmarks for evaluating RAG performance, as creating high-quality datasets and experimenting on them entail significant costs. These benchmarks can be classified into two types: reference-required and reference-free evaluation. Reference-free evaluation frameworks, such as RAGAS (Es et al., 2023) and ARES (Saad-Falcon et al., 2023), use LLM-generated data to evaluate RAG systems on contextual relevance, faithfulness, and informativeness. These frameworks do not depend on ground truth references, but only assess the coherence of the generated text with the retrieved context. This approach may be unreliable if the retrieved external information is low-quality.
Consequently, reference-required evaluations remain the predominant method for assessing RAG systems. Existing benchmarks for reference-required evaluations, such as RGB (Chen et al., 2023b) and NQ (Kwiatkowski et al., 2019). do have their limitations. First, they all rely on question answering tasks to measure the performance of RAG systems. Question answering is not the only RAG application scenario, and an optimization strategy that works well for question answering may not be generalized to other scenarios. Thus, these benchmarks may not capture the full potential of RAG systems. Second, in the experiments, current evaluations usually focus on evaluating the LLM part of the RAG pipeline, while ignoring the retrieval model and external knowledge base construction. These components are also crucial for RAG systems, as they determine the quality and relevance of retrieved passages and external knowledge. Therefore, a comprehensive evaluation of the RAG system may not be obtained using any existing benchmarks.
To evaluate the performance of RAG in different application scenarios, we need a comprehensive benchmark that covers more than just the question answering task.  Lewis et al. (2020) argue that the core of RAG systems is their interactive way of combining LLMs with external knowledge sources. And following  (Kilov, 1990), we can group any interaction with external knowledge sources into four basic actions: create, read, update, and delete, which are also known as CRUD actions (Truica et al., 2015). Therefore, we can use the CRUD framework to classify the RAG systems’ application scenarios. As shown in Figure 1, each CRUD category demonstrates different capabilities of the RAG system:
@@
  • In ”READ”, the system uses external knowledge retrieval to answer questions, solving problems in question-answering, dialogue, and reasoning, and increasing understanding of the input text.
  • In ”UPDATE”, the system fixes errors in the input text using retrieved content, correcting spelling, grammar or factual errors to make the text better.
@@
To evaluate the RAG system in these four scenarios, we introduce CRUD-RAG, a comprehensive, large-scale Chinese RAG benchmark. CRUD-RAG consists of four evaluation tasks: text continuation, question answering (with single-document and multi-document questions), hallucination modification, and open-domain multi-document summarization, which respectively correspond to the CRUD-RAG classification of RAG application scenarios. We construct CRUD-RAG by crawling the latest high-quality news data from major news websites in China, which were not exposed to the LLMs during training. We then automatically create datasets using GPT-4 based on these news data. For the multi-document summarization task, we apply a reverse construction strategy. We first generate news events and their summaries using GPT-4. Then, we use these events as keywords to search for 10 related and non-duplicate reports from the web, which we add to our retrieval database. During evaluation, the RAG system will use the retrieval database to generate summaries for the events. For the text continuation task, we split the news text into a beginning and a continuation paragraph. We then use each sentence in the continuation paragraph as a keyword to search for 10 related reports on the Web. We remove any duplicate content and add the reports to the retrieval database. For the single-document QA task, we use the RGB (Chen et al., 2023b) construction method. For the multi-document QA task, we use the Chain-of-Thought technology to help the model identify common and different aspects among documents, and then generate questions based on these aspects with increasing difficulty. For the hallucination modification task, we use the annotations in the UHGEval dataset and correct hallucinations with GPT-4. We also include the real news in UHGEval in the retrieval database.
@@
  • High-quality evaluation datasets: We constructed diverse datasets for different evaluation tasks, based on the application scenarios of RAG. These tasks include text continuation, multi-document summarization, question answering and hallucination modification.
  • Extensive experiments: we performed extensive experiments on our own benchmark, using various metrics to measure the performance of RAG systems. Based on our experiments, we offered useful guidance for future researchers and RAG system developers.
@@ 2. Related Work
2.1. RAG: A Neural Text Generation Framework that Combines Retrieval and Generation
LLMs excel in text generation but also confront challenges such as outdated knowledge and the generation of hallucinatory content (He et al., 2022; Cao et al., 2020; Raunak et al., 2021). In response to these challenges, RAG, also referred to as RALM (Retrieval-Augmented Language Models), incorporate external knowledge to generate responses characterized by enhanced accuracy and realism (Shi et al., 2023). This is particularly critical in domains that heavily depend on precision and reliability, including but not limited to the legal, medical, and financial sectors. Retrieval models have been promoting the development of language models (Liu et al., 2024; Zhai and Lafferty, 2004; Ganguly et al., 2015).
@@ 2.1. RAG: A Neural Text Generation Framework that Combines Retrieval and Generation
2.2. RAG Benchmark: Evaluation Datasets and Methods for Retrieval-Augmented Generation
@@
When investigating the development and optimization of RAG, the effective evaluation of their performance becomes a fundamental concern. LangChain provides benchmark tasks, such as LangChain Docs Q&A and Semi-structured Reports (Langchain, 2023), designed to assess various RAG architectures. These datasets are constructed from snapshots of Python documentation and PDFs containing tables and charts. They emphasize the model’s capability to handle structured and semi-structured data. Evaluation standards encompass the accuracy of answers and the faithfulness of model responses. Utilizing large models for question-answering generation has emerged as a prevalent approach in building evaluation datasets. For instance, RGB (Chen et al., 2023b) creates its evaluation dataset by gathering recent news reports and employing LLM to generate relevant events, questions, and answers. Conversely, ARES (Saad-Falcon et al., 2023). relies on generating synthetic queries and answers, leveraging the FLAN-T5 XXL model. These methods not only showcase the RAG system’s proficiency in handling real-time data but also illustrate the utility of automation and synthetic data in the evaluation process. For evaluating the capabilities of models across various professional domains, the Instruct-Benchmark-Tester dataset encompasses a range of question types, with a particular focus on financial services, legal, and intricate business scenarios (Oberst, 2023).
@@
Reference-free evaluation methods, including TruLens-Eval (Ferrara et al., 2024), RAGAS (Es et al., 2023), and ARES (Saad-Falcon et al., 2023), provide distinct viewpoints for evaluating the performance of RAG systems, particularly concerning context relevance, answer faithfulness, and answer relevance. TruLens-Eval (Ferrara et al., 2024) introduces the RAG Triad as an innovative approach to evaluate hallucination issues within the RAG architecture, encompassing context relevance, groundedness, and answer relevance. RAGAS (Es et al., 2023), serving as a reference-free evaluation framework, concentrates on assessing the retrieval system’s capacity to identify pertinent and concentrated context passages, along with the LLMs’ proficiency in faithfully and accurately leveraging these passages. In contrast to RAGAS, which depends on a predefined set of heuristically crafted prompts, ARES generates tailored LLMs judges for each aspect of a RAG pipeline, leading to a substantial enhancement in evaluation precision and accuracy when compared to existing methods such as RAGAS. Furthermore, ARES (Saad-Falcon et al., 2023) employs prediction-powered inference to offer statistical assurances for its scoring, generating confidence intervals. ARES emphasizes three evaluation scores: context relevance, answer faithfulness, and answer relevance, highlighting the importance of a proficient RAG system in identifying relevant contexts and producing both faithful and relevant answers. Regarding evaluation methods,  (Liu et al., 2023) places an emphasis on assessing the credibility and accuracy of responses generated by generative search engines through manual inspection. Nonetheless, manual evaluation possesses drawbacks, including high costs and challenges in scalability. Hence, rule-based evaluation metrics such as accuracy, exact match, rouge, or self-devised metrics like rejection rate, error detection rate, and correction rate continue to be widely adopted in the field. Furthermore, employing large language models for evaluation closely approximates manual evaluation outcomes.
@@ 3. CRUD-RAG: A Comprehensive Chinese Benchmark for RAG
However, The current RAG benchmarks have several drawbacks: they only evaluate question answering tasks (Qu et al., 2020; Aliannejadi et al., 2019; Yu et al., 2020), ignoring other diverse application of RAG. The optimization strategy for question answering tasks may not suit other tasks; And in the evaluation experiment, current RAG benchmarks only account for the large language model component in the RAG pipeline, disregarding the vital roles of retrieval database construction and retrieval.
To address the shortcomings of previous benchmarks, we introduce CRUD-RAG, a comprehensive Chinese benchmark for RAG. Figure 2 illustrates the features of our CRUD-RAG benchmark. It classifies the RAG application scenarios into four categories: Create, Read, Update, Delete, and constructs appropriate evaluation tasks and datasets for each category. Besides, in the experiments, we will assess the impact of various components of RAG, such as chunk size, retrieval strategy, top-k, LLM, etc., on all tasks.
In the following section, we will describe the evaluation tasks and the datasets that we design for each RAG application scenario type. Table 2 presents the size and composition of our datasets, and Figure 3 illustrates an example of our datasets.
@@
Dataset Name
Dataset Size
Components
Evaluation objectives
Text Continuation
10,728
An initial part of an article, followed by its extension or completion.
Evaluate the RAG system’s performance in ”Create” scenarios(Creative generation).
@@
3,199
A collection of question-answer pairs, where the answer is directly extractable from a document passage.
Evaluate the RAG system’s performance in ”Read” scenarios(Knowledge-intensive application).
@@
3,192
A collection of question-answer pairs, where the answer requires synthesis of information from 2 different document sources.
The objective is the same as 1-document QA, but it also examines the reasoning ability of combining 2 documents.
@@
3,189
A collection of question-answer pairs, where the answer requires synthesis of information from 3 different document sources.
The objective is the same as 1-document QA, but it also examines the reasoning ability of combining 3 documents.
@@
5,130
Some sentences containing errors, and the sentence with the errors fixed.
Evaluate the RAG system’s performance in ”Update” scenarios(Error correction application).
@@
10,728
A one-sentence headline of an article, followed by a brief summary of the article.
Evaluate the RAG system’s performance in ”Delete” scenarios(Summarization).
Retrieval Database
86,834
As the knowledge base for the RAG system, we expect the RAG system to retrieve relevant content from the knowledge base to address the above tasks.

### v3

@@ Abstract.
Retrieval-Augmented Generation (RAG) is a technique that enhances the capabilities of large language models (LLMs) by incorporating external knowledge sources. This method addresses common LLM limitations, including outdated information and the tendency to produce inaccurate ”hallucinated” content. However, evaluating RAG systems is a challenge. Most benchmarks focus primarily on question answering applications, neglecting other potential scenarios where RAG could be beneficial. Accordingly, in the experiments, these benchmarks often assess only the LLM components of the RAG pipeline or the retriever in knowledge-intensive scenarios, overlooking the impact of external knowledge base construction and the retrieval component on the entire RAG pipeline in non-knowledge-intensive scenarios. To address these issues, this paper constructs a large-scale and more comprehensive benchmark, and evaluates all the components of RAG systems in various RAG application scenarios. Specifically, we refer to the CRUD actions that describe interactions between users and knowledge bases, and also categorize the range of RAG applications into four distinct types—Create, Read, Update, and Delete (CRUD). ”Create” refers to scenarios requiring the generation of original, varied content. ”Read” involves responding to intricate questions in knowledge-intensive situations. ”Update” focuses on revising and rectifying inaccuracies or inconsistencies in pre-existing texts. ”Delete” pertains to the task of summarizing extensive texts into more concise forms. For each of these CRUD categories, we have developed different datasets to evaluate the performance of RAG systems. We also analyze the effects of various components of the RAG system, such as the retriever, context length, knowledge base construction, and LLM. Finally, we provide useful insights for optimizing the RAG technology for different scenarios [^1] The source code is available at GitHub: https://github.com/IAAR-Shanghai/CRUD_RAG.
@@ 1. Introduction
Although the effectiveness of retrieval-augmented strategies has been proven through extensive practice, their implementation still requires a significant amount of tuning. The overall performance of the RAG system is affected by multiple factors, such as the retrieval model, construction of the external knowledge base, and language model. Therefore, automatic evaluation of RAG systems is crucial. Currently, there are only a few existing benchmarks for evaluating RAG performance, as creating high-quality datasets and experimenting with them entail significant costs. These benchmarks can be classified into two types: reference-required and reference-free evaluation. Reference-free evaluation frameworks, such as RAGAS (Es et al., 2023) and ARES (Saad-Falcon et al., 2023), use LLM-generated data to evaluate RAG systems on contextual relevance, faithfulness, and informativeness. These frameworks do not depend on ground truth references, but only assess the coherence of the generated text with the retrieved context. This approach may be unreliable if the retrieved external information is low-quality.
Consequently, reference-required evaluations remain the predominant method for assessing RAG systems. Existing benchmarks for reference-required evaluations, such as RGB (Chen et al., 2023b) and NQ (Kwiatkowski et al., 2019). do have their limitations. First, they all rely on question answering tasks to measure the performance of RAG systems. Question answering is not the only RAG application scenario, and an optimization strategy that works well for question answering may not be generalized to other scenarios. Thus, these benchmarks may not capture the full potential of RAG systems. Second, in the experiments, current evaluations usually focus on evaluating the LLM part of the RAG pipeline, or focus on retriever performance in the knowledge-intensive scenario (Petroni et al., 2020), while ignoring the retrieval methods in non-knowledge-intensive scenarios and external knowledge base construction. These components are also crucial for RAG systems. Therefore, a comprehensive evaluation of the RAG system may not be obtained using any existing benchmarks.
To evaluate the performance of RAG in different application scenarios, we need a comprehensive benchmark that covers more than just the question-answering task.  Lewis et al. (2020) argue that the core of RAG systems is their interactive way of combining LLMs with external knowledge sources. And following  (Kilov, 1990), we can group any interaction with external knowledge sources into four basic actions: create, read, update, and delete, which are also known as CRUD actions (Truica et al., 2015). Therefore, we can use the CRUD framework to classify the RAG systems’ application scenarios. As shown in Figure 1, each CRUD category demonstrates different capabilities of the RAG system:
@@
  • In ”READ”, the system uses external knowledge retrieval to answer questions, solve problems in question-answering, dialogue, and reasoning, and increase understanding of the input text.
  • In ”UPDATE”, the system fixes errors in the input text using retrieved content, correcting spelling, grammar, or factual errors to make the text better.
@@
To evaluate the RAG system in these four scenarios, we introduce CRUD-RAG, a comprehensive, large-scale Chinese RAG benchmark. CRUD-RAG consists of four evaluation tasks: text continuation, question answering (with single-document and multi-document questions), hallucination modification, and open-domain multi-document summarization, which respectively correspond to the CRUD-RAG classification of RAG application scenarios. We construct CRUD-RAG by crawling the latest high-quality news data from major news websites in China, which aims to minimize the likelihood of LLMs encountering these data during training. Then, we automatically create datasets using GPT-4 based on these news data. For the multi-document summarization task, we apply a reverse construction strategy. We first generate news events and their summaries using GPT-4. Then, we use these events as keywords to search for 10 related and non-duplicate reports from the web, which we add to our retrieval database. During evaluation, the RAG system will use the retrieval database to generate summaries for the events. For the text continuation task, we split the news text into a beginning and a continuation paragraph. We then use each sentence in the continuation paragraph as a keyword to search for 10 related reports on the Web. We remove any duplicate content and add the reports to the retrieval database. For the single-document QA task, we use the RGB (Chen et al., 2023b) construction method. For the multi-document QA task, we use the Chain-of-Thought technology to help the model identify common and different aspects among documents, and then generate questions based on these aspects with increasing difficulty. For the hallucination modification task, we use the annotations in the UHGEval dataset and correct hallucinations with GPT-4. We also include the real news in UHGEval in the retrieval database.
@@
  • High-quality evaluation datasets: We constructed diverse datasets for different evaluation tasks, based on the application scenarios of RAG. These tasks include text continuation, multi-document summarization, question answering, and hallucination modification.
  • Extensive experiments: we performed extensive experiments on our benchmark, using various metrics to measure the performance of RAG systems. Based on our experiments, we offered useful guidance for future researchers and RAG system developers.
@@ 2. Related Work
2.1. Retrieval-Augmented Generation
LLMs excel in text generation but also confront challenges such as outdated knowledge and the generation of hallucinatory content (He et al., 2022; Cao et al., 2020; Raunak et al., 2021). In response to these challenges, RAG, also referred to as RALM (Retrieval-Augmented Language Models), incorporates external knowledge to generate responses characterized by enhanced accuracy and realism (Shi et al., 2023). This is particularly critical in domains that heavily depend on precision and reliability, including but not limited to the legal, medical, and financial sectors. Retrieval models have been promoting the development of language models (Liu et al., 2024; Zhai and Lafferty, 2004; Ganguly et al., 2015).
@@ 2.1. Retrieval-Augmented Generation
2.2. RAG Benchmarks
@@
Citation
Yes EN
 (Gao et al., 2023c; Ji et al., 2024)
Questions, answers
and contexts
(with citations)
Fluency, Correctness,
Citation quality
Self-devised metrics,
Human evaluation
Citation
Yes EN
 (Xu et al., 2024)
Questions, answers
and contexts
(with citations)
1948
Location citation recall, Location precision,
The coefficient of variation of citation locations
Self-devised metrics
Citation
Yes EN
 (Li et al., 2024)
Questions, answers
and contexts
(with citations)
3422
Accuracy of factual and non-factual,
AUC-PR, and so on
Self-devised metrics,
Common metrics
Citation
Yes EN
 (Zhang et al., 2024)
Statement-citation Pairs
12681
Correlation, Classification performance,
Retrieval effectiveness, Faithfulness
Self-devised metrics,
Common metrics,
Human evaluation
Citation
Yes EN
 (Cao and Wang, 2024)
Paragraphs
(with citations)
10000
Citation density,
The coverage of reference facts
Self-devised metrics
@@
When investigating the development and optimization of RAG, the effective evaluation of their performance becomes a fundamental concern. Table 1 shows some commonly used benchmarks for evaluating RAG. LangChain provides benchmark tasks, such as LangChain Docs Q&A and Semi-structured Reports (Langchain, 2023), designed to assess various RAG architectures. These datasets are constructed from snapshots of Python documentation and PDFs containing tables and charts. They emphasize the model’s capability to handle structured and semi-structured data. Evaluation standards encompass the accuracy of answers and the faithfulness of model responses. Utilizing large models for question-answering generation has emerged as a prevalent approach in building evaluation datasets. For instance, RGB (Chen et al., 2023b) creates its evaluation dataset by gathering recent news reports and employing LLM to generate relevant events, questions, and answers. Conversely, ARES (Saad-Falcon et al., 2023). relies on generating synthetic queries and answers, leveraging the FLAN-T5 XXL model. These methods not only showcase the RAG system’s proficiency in handling real-time data but also illustrate the utility of automation and synthetic data in the evaluation process. For evaluating the capabilities of models across various professional domains, the Instruct-Benchmark-Tester dataset encompasses a range of question types, with a particular focus on financial services, legal, and intricate business scenarios (Oberst, 2023).
@@
Reference-free evaluation methods, including TruLens-Eval (Ferrara et al., 2024), RAGAS (Es et al., 2023), and ARES (Saad-Falcon et al., 2023), provide distinct viewpoints for evaluating the performance of RAG systems, particularly concerning context relevance, answer faithfulness, and answer relevance. TruLens-Eval (Ferrara et al., 2024) introduces the RAG Triad as an innovative approach to evaluate hallucination issues within the RAG architecture, encompassing context relevance, groundedness, and answer relevance. RAGAS (Es et al., 2023), serving as a reference-free evaluation framework, concentrates on assessing the retrieval system’s capacity to identify pertinent and concentrated context passages, along with the LLMs’ proficiency in faithfully and accurately leveraging these passages. In contrast to RAGAS, which depends on a predefined set of heuristically crafted prompts, ARES generates tailored LLMs judges for each aspect of a RAG pipeline, leading to a substantial enhancement in evaluation precision and accuracy when compared to existing methods such as RAGAS. Furthermore, ARES (Saad-Falcon et al., 2023) employs prediction-powered inference to offer statistical assurances for its scoring, generating confidence intervals. ARES emphasizes three evaluation scores: context relevance, answer faithfulness, and answer relevance, highlighting the importance of a proficient RAG system in identifying relevant contexts and producing both faithful and relevant answers. Regarding evaluation methods,  (Liu et al., 2023) places an emphasis on assessing the credibility and accuracy of responses generated by generative search engines through manual inspection. Nonetheless, manual evaluation possesses drawbacks, including high costs and challenges in scalability. Hence, rule-based evaluation metrics such as accuracy, exact match, rouge, or self-devised metrics like rejection rate, error detection rate, and correction rate continue to be widely adopted in the field. Furthermore, employing LLMs for evaluation closely approximates manual evaluation outcomes.
2.3. Citation-Enhanced RAG
In traditional RAG methods, despite the rich information sources provided by retrieved contexts for text generation, these models often do not explicitly require responses to provide corresponding citations, making traceability difficult. Therefore, enhancing text verifiability by introducing citation links, i.e., explicit references, has become an important research direction in the RAG field (Gao et al., 2023c; Xu et al., 2024).
Providing citation indicators in the response text offers several clear benefits. First, users can easily verify the claims made by LLMs based on the provided citations, thus improving the transparency and credibility of the text. Second, if the text generated by LLMs adheres faithfully to the cited contexts, it can significantly improve its accuracy and reduce the phenomenon of ”hallucinations” (Gao et al., 2023c). Given this, generating high-quality citations and evaluating the quality of citation generation have become crucial elements of assessing RAG performance. Constructing appropriate prompts directly through the retrieval context to guide the model in generating corresponding citations constitutes a direct and effective method of citation generation (Ji et al., 2024).
In terms of evaluation, early research primarily focused on the fluency, accuracy, and basic citation quality of the text generated by LLMs (Gao et al., 2023c; Li et al., 2024). For example, Rashkin et al. proposed the ”Attributable to Identified Sources” (AIS) score (Rashkin et al., 2023), which serves as a valuable tool for measuring the degree to which generated text is faithful to its sources. As research progressed, scholars recognized the need for more detailed evaluation methods to differentiate between various levels of citation support. By creating specialized datasets such as SCIFI (Cao and Wang, 2024), researchers can more precisely evaluate fine-grained citations at the clause level in texts generated by LLMs. The ALiiCE framework (Xu et al., 2024), by analyzing the atomic structure of sentence claims, introduced fine-grained evaluation metrics, such as location citation recall and precision, and the coefficient of variation of citation locations, to more granularly evaluate the quality of citation generation in RAG (Xu et al., 2024). In practical applications,  (Zhang et al., 2024) found that RAG requires more complex evaluation frameworks to distinguish between various levels of citation support by comparing different fidelity metrics. These RAG evaluation methods not only consider the presence of citations but also their accuracy and relevance.
While Citation-Enhanced RAG delves deeply into the specific domain of citation generation, aiming to improve the credibility and accuracy of text generated by RAG systems, our benchmark provides a comprehensive evaluation framework encompassing various aspects of RAG systems and multiple application scenarios.
@@ 3. CRUD-RAG: A Comprehensive Chinese Benchmark for RAG
However, The current RAG benchmarks have several drawbacks: they only evaluate question answering tasks (Qu et al., 2020; Aliannejadi et al., 2019; Yu et al., 2020), ignoring other diverse application of RAG. The optimization strategy for question-answering tasks may not suit other tasks; And in the evaluation experiment, current RAG benchmarks only account for the LLM component in the RAG pipeline, or the retriever in the knowledge-intensive scenario, disregarding the vital roles of retrieval database construction and retrieval in non-knowledge-intensive scenarios.
To address the shortcomings of previous benchmarks, we introduce CRUD-RAG, a comprehensive Chinese benchmark for RAG. Figure 2 illustrates the features of our CRUD-RAG benchmark. It classifies the RAG application scenarios into four categories: Create, Read, Update, and Delete, then we construct appropriate evaluation tasks and datasets for each category. Besides, in the experiments, we will assess the impact of various components of RAG, such as chunk size, retrieval strategy, top-k, LLM, etc., on all tasks.
In the following section, we will describe the evaluation tasks and the datasets that we design for each RAG application scenario type. We select text continuation, question answering (single and multi-document), hallucination modification, and multi-document summarization as representative tasks in the CRUD (Create, Read, Update, Delete) scenario and construct corresponding datasets. The summarization (D) and continuation (C) datasets were constructed simultaneously, since the construction of both datasets requires the use of a search engine. They will be discussed together in the following section. The construction of the question-answering(R) and hallucination modification(D) datasets is relatively independent. To maintain narrative coherence, we will introduce the dataset construction process in the order of DCRU. Table 2 presents the size and composition of our datasets, and Figure 3 illustrates an example of our datasets.
@@
Dataset Name Dataset Size Components Evaluation objectives
Text Continuation 10,728 An initial part of an article, followed by its extension or completion. Evaluate the RAG system’s performance in ”Create” scenarios(Creative generation).
@@
3,199 A collection of question-answer pairs, where the answer is directly extractable from a document passage. Evaluate the RAG system’s performance in ”Read” scenarios(Knowledge-intensive application).
@@
3,192 A collection of question-answer pairs, where the answer requires synthesis of information from 2 different document sources. The objective is the same as 1-document QA, but it also examines the reasoning ability of combining 2 documents.
@@
3,189 A collection of question-answer pairs, where the answer requires synthesis of information from 3 different document sources. The objective is the same as 1-document QA, but it also examines the reasoning ability of combining 3 documents.
@@
5,130 Some sentences containing errors, and the sentence with the errors fixed. Evaluate the RAG system’s performance in ”Update” scenarios(Error correction application).
@@
10,728 A one-sentence headline of an article, followed by a brief summary of the article. Evaluate the RAG system’s performance in ”Delete” scenarios(Summarization).
Retrieval Database 86,834 As the knowledge base for the RAG system, we expect the RAG system to retrieve relevant content from the knowledge base to address the above tasks. ———